The element Kr has van der Waals constants a = 2.325 atm•L2/mol2 and b = 0.0396 L/mol. Using both the ideal gas law and van der Waals’s equation, calculate the pressure expected for 30 mol of Kr gas in a 6.00-L container at 25 °C.
Answer #1

1) Using ideal gas law

Given:

V = 6.0 L

n = 30 mol

T = 25.0 oC

= (25.0+273) K

= 298 K

use:

P * V = n*R*T

P * 6 L = 30 mol* 0.08206 atm.L/mol.K * 298 K

P = 122 atm

Answer: 122 atm

2) suing van der Waals equation

Given:

V = 6.0 L

n = 30.0 mol

R = 0.08206 atm.L/mol.K

T = 298.0 K

a = 2.325 atm.L^2/mol^2

b = 0.0396 L/mol

use:

(P+an^2/V^2)*(V-nb) = n*R*T

(P + 2.325*30.0^2/6.0^2)*(6.0-30.0*0.0396) = 30.0*0.08206*298.0

(P + 58.125)*(4.812) = 733.6164

P + 58.125 = 152.4556

P = 94.3 atm

Answer: 94.3 atm